# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v2

@@ 3.3.1 Context-Aware Attack Activation
In our hazard analysis, we define the accident as the adverse event of forward collision with the lead vehicle (measured by a zero or negative relative distance between the front object and the Ego vehicle). This kind of accident can happen as a result of the Ego vehicle transitioning into a hazardous state that violates the safe following distance with the lead vehicle.
[…]
TABLE II: Partial safety context table for an ACC system.
@@ 4 RS>0 No
3.3.2 Adaptive Adversarial Patch Generation
@@
Optimization-based Adversarial Patch Generation. To address challenge C2, we formulate the attack as the following optimization problem:
Figure 3: Optimization-based adversarial patch generation.
min $\sum_{d\in{RD}_{t}}{-\triangledown g(d,\theta)}+\lambda||X^{\prime}_{t}||_{p}$ (1)
s.t. $X^{\prime}_{t}=X_{t}+Patch$ (2)
$[{RD},{V_{Lead}}]_{t}=LVD_{\theta}(X^{\prime}_{t-1})$ (3)
$u_{t}=ACC(s_{t},[{RD},{V_{Lead}}]_{t})$ (4)
@@
$Patch=\Delta*M$ (6)
$Patch\in[\mu-\sigma,\mu+\sigma]$ (7)
$Area(Patch)\subset BBox(V_{lead})$ (8)
where Eq. 1 is an objective function that aims at accelerating the Ego vehicle as soon as possible to cause a forward collision with the lead vehicle. Directly decreasing the probability of a lead vehicle or its bounding box (BBox) cannot change the ACC system behavior. We instead design an objective function that increases the relative distance between the lead vehicle and the Ego vehicle as much as possible while keeping the perturbation value of the adversarial patch imperceptible to human eyes.
@@
In Eq. 1, $g(d)$ is an approximate polynomial function of $d$ that fits the trend of the trajectory of the relative distance ${RD}_{t}$ between the lead vehicle and Ego vehicle, predicted by the DNN model with weight parameters $\theta$. "-" is a negative sign that converts our goal of maximizing the relative distance to minimizing the proposed objective function. For example, when the gradient of the relative distance trajectory, $g(d)$, is negative, minimizing "$-\triangledown g(d)$" will slow down the decrease of $g(d)$ and is equivalent to maximizing the relative distance. We adopt the gradient of $g(d)$ in the objective function instead of using ${RD}_{t}$ itself in order to avoid sharp changes in the predicted relative distance value, which might be easily detected by some anomaly detection mechanisms.
In Eq. 2, the perturbation is added to the original image input $X_{t}\in\mathbb{R}^{H\times W\times C}$ in the form of an adversarial patch $Patch\in\mathbb{R}^{H\times W\times C}$, represented as a matrix of pixels with height H, width W, and C color channels. $\lambda$ is the weight parameter of the p-norm regularization term, designed to minimize the perturbation value of the patch for stealthiness. We limit the perturbation value within the Kalman filter noise parameters ($\mu$,$\sigma$) (Eq. 7), which ensures the perturbation is not corrected by the sensor fusion. We also constrain the adversarial patch inside the BBox of the lead vehicle (Eq. 8) to enhance attack effectiveness, minimize the perturbation area for stealthiness, and reduce computational cost.
@@
Primary Attribution Detection and Patch Update. As mentioned in Section 3.2, a major challenge (C3) in the design of runtime attacks is the changes in the size and location of the lead vehicle in the perceived image frames. To address this challenge, the attacker needs to update the generated patch dynamically according to the approximate DNN outputs. In this paper, we adopt an object detection method [50] to detect and track the real-time location and size of the lead vehicle and concentrate the attack perturbation within the detected BBox of the lead vehicle. In production ACC with object detection features [22], the attack can skip this step and use the stock prediction result.
@@
After getting the BBox, we utilize a primary attribution algorithm [48] to quantify the relationship between input features and output predictions. Through this exploration, we try to identify the important pixels within the BBox of the lead vehicle that contribute the most to the predictions of relative distance. The input pixels with high weights identified by the attribution algorithm are marked by unit value in the mask matrix $M$, and the remaining pixels are assigned zero values. This mask matrix is then multiplied by the perturbation $\Delta$ to generate the adversarial $Patch$ (in Eq. 6). This step is useful as it can filter out non-important pixels in the inputs to reduce the number of perturbed pixels to improve the effectiveness of optimization-based attacks and reduce computation cost for runtime attacks.
@@
Finally, we develop a new initialization algorithm to shift the patch position and adjust its size when the detected BBox changes (Eq. 9-13). We shift the attack vector toward the new position of the detected BBox of the lead vehicle with a magnitude of ($x_{1}-x_{0}$,$y_{1}-y_{0}$), where ($x_{0}$,$y_{0}$) and ($x_{1}$,$y_{1}$) are the centers of the previous BBox and current BBox, respectively (Eq. 9). We then expand the adversarial patch attack vector ($Patch$) to the dimension that matches the size of the newly detected BBox of the lead vehicle. Instead of reinitializing the whole attack vector matrix with random or zero values, which will reset the whole optimization process, we keep the previous patch values and intermediate variables and only initialize the newly expended units (Eq. 12-13). An example is shown in Fig. 5.
Fig. 5 shows a visualization of how the adversarial patch affects the DNN predictions.
4 Safety Intervention Simulation
@@
To evaluate the safety of DNN-based ACC systems under attacks, we enhance the default OpenPilot and CARLA simulation platform (see Section 2.3) to be more representative of real-world ADAS, by developing a safety intervention simulator and mechanisms for priority-based dispatching of control commands to CARLA and fusion of camera and radar data (see Appendix A). An overview of the simulation platform is shown in Fig. 6 (with the orange parts representing our new implementations) and presented next.
@@
To fill the gap in considering safety interventions and address the challenge of ensuring the combination of human driver and vehicle safe (see Section 2.2), we implement and integrate three levels of safety interventions in the OpenPilot software (see Fig. 6), including ADAS safety features (e.g., AEB and FCW), basic car safety constraint checking on control commands, and driver interventions.
@@
The AEBS processes inputs from LVD outputs (after sensor fusion), including relative distance ($RD$) and relative speed ($RV$), and current speed $V_{Ego}$ (see Fig. 1). The average driver reaction time ($T_{react}$) is set to the commonly used constant value of 2.5s [17, 11]. Time thresholds, namely $ttc$ (time to collision), $t_{fcw}$ (FCW time), $t_{pb1}$ (1st phase partial brake time), $t_{pb2}$ (second phase partial brake time), and $t_{fb}$ (full brake time), are then calculated as follows:
$ttc$ $=RD/RV$ (14)
@@
As shown int Fig. 8, when $ttc$ falls below $t_{fcw}$, $t_{pb1}$, $t_{pb1}$, and $t_{pb1}$, a corresponding action (warning or brake with 90%, 95%, 100% force) is executed. Applying the brake value blocks other controls from the ADAS. More details of AEBS design and testing are provided in Appendix C.
In reality, when OpenPilot is installed on a car model, some models lose the AEBS functionality [31], while others retain it. Also, AEBS might rely on a separate ADAS camera [55], distinct from the OpenPilot camera, and a potential compromise of the AEBS camera data is possible. Thus, we consider three scenarios for AEBS interventions: (1) AEBS enabled, and AEBS camera data is uncompromised; (2) AEBS enabled, but AEBS camera data is compromised; and (3) AEBS is disabled (see Section 5.3.2 and Table VI).
Safety Constraint Checker. The OpenPilot safety mechanisms are implemented in its control software and the Panda CAN interface. Panda is a universal OBD adapter developed by Comma.ai [56] that provides access to almost all car sensors through the CAN bus and also enforces some safety constraints over output commands. However, when integrated with the CARLA driving simulator, OpenPilot does not utilize Panda software and hardware and, thus, Panda safety checks are inactive.
To be as realistic as the actual OpenPilot on the road, we add a virtual Panda module in our simulation that copies the exact logic of Panda software [56]. Specifically, as shown in Fig. 6, the virtual Panda decodes the CAN packages sent by the ACC and checks whether their checksum is correct and the control command values are within predefined thresholds [56]. For example, to ensure safety, the maximum acceleration and deceleration of the vehicle shall be limited between $m/s^{2}$ and -3.5 $m/s^{2}$, respectively [27]. Only the commands that pass the Panda safety checks are sent to the simulated vehicle actuators. In CARLA simulator, the final control commands are truncated within the range of [0,1].
Driver Reaction Simulator. To assess driver interventions, we develop a driver reaction simulator. The simulated driver is notified when any safety alerts are raised by the ADAS (e.g., FCW) or when the driver observes any abnormalities in the vehicle’s status or camera user interface (UI). These ACC abnormalities include hard braking, unexpected acceleration, the vehicle’s speed exceeding the cruising speed by more than 10%, or the mean perturbation value in the camera UI surpassing a noticeable threshold, with a default value of 15% representing an alert driver ($Patch.mean()>0.15$). We assume a very alert driver who can notice any anomalies that occur within a single control cycle (10ms). A predefined emergency response will be issued by the driver accordingly (see Table III), taking effect 2.5 seconds later (average driver reaction time). To mimic the human driver’s braking behavior, we adopt an approximate braking curve function from prior work [17].
TABLE III: Driver simulator: activation conditions and reactions.
@@
Priority-based Control Command Dispatcher. With multiple safety mechanisms in place, there might be conflicts among the control commands issued by the OpenPilot ACC controller and those generated by the safety interventions. To resolve such conflicts, we design a command dispatcher to transmit output control commands to the CARLA actuators from various sources (e.g., ACC, AEB, simulated driver) based on their priorities, with high-priority commands overwriting low-priority ones (see Fig. 6 and Fig. 8). The simulated driver’s actions have a higher priority than regular ACC outputs, and control actions from the AEB have the highest priority. The human driver’s actions will be executed 2.5s (average driver reaction time) after safety alerts (e.g., FCW) or noticing other ACC malfunctions (see Table III). ACC commands will be blocked or disengaged when AEB or driver interventions are triggered.
Figure 7: AEBS.
@@
5 Evaluation in Simulated Environment
This section presents the evaluation of our proposed context-aware attack strategy (referred to as CA-Opt) using the enhanced simulation platform presented in Section 4.
5.1 Methodology
We study the following research questions by comparing the effectiveness of CA-Opt attack to several baseline attack methods in causing safety hazards (Section 5.2) and evading different safety interventions (Section 5.3):
@@ 5.1 Methodology
Baselines. We design three baseline attack strategies to answer these questions (see Table IV).
TABLE IV: Overview of proposed and baseline attack strategies.
@@
CA-FGSM Context-Aware Context-Aware FGSM 1000
@@
To assess the effectiveness of our optimization-based adversarial patch method in strategically selecting the attack values, we compare it to two baselines: CA-Random, which introduces random perturbations to perception inputs, and CA-FGSM, which uses a fast gradient sign method [21] to determine the perturbation values. Both baselines use the same context-aware method as the proposed CA-Opt attack strategy to choose the attack start time and duration. To evaluate our method’s efficiency in selecting the timing and duration of attacks, we design another baseline (Random-Opt) that selects a random start time uniformly distributed within [5, 40] seconds and a random attack duration uniformly distributed within [0.5, 2.5] seconds. To concentrate only on the effect of different start times and durations, Random-Opt shares the same adversarial patch generation method as CA-Opt. For a fair comparison, we confine perturbations to the detected bounding box (BBox) of the lead vehicle and update the BBox size and position with our proposed patch updating algorithm (Section 5).
@@
Our experiments are done on Ubuntu 20.04 LTS, with OpenPilot v0.8.9 and CARLA v9.11. A single simulation of OpenPilot contains 5000 time steps, and each step lasts about 10 ms, which equals 50 seconds in total. For Random-Opt, we randomly select ten start times and five durations for each test scenario and repeat them five times, which results in 1,000 simulations. The same total number of simulations is done for CA-Opt and other baselines.
5.2 Attack Success Rate in Causing Hazards

### v3

@@ III-C1 Context-Aware Attack Activation
In our hazard analysis, we define the accident as the adverse event of forward collision (measured by a zero or negative relative distance between the Ego vehicle and the lead vehicle or a front object). This can happen as a result of the Ego vehicle transitioning into a hazardous state that violates the safe following distance with the lead vehicle.
To determine the critical system contexts, we assess all the combinations of system states and ACC control actions (e.g., acceleration, deceleration) to identify the specific combinations that are most likely to lead to hazards. Table III shows part of the safety context table for ACC with a focus on the acceleration commands. Here, the critical system context is described as when the Headway Time (HWT, the time the Ego vehicle takes to drive the relative distance ($RD$) from the lead vehicle with the current speed) is less than a safety limit HWTsafe (e.g., 2-3s), and the Ego vehicle is faster than the lead vehicle ($V_{Ego}>V_{Lead}$). Under such system context, an acceleration command induced by the camera perception attacks or other reasons will most likely lead to a forward collision hazard. This high-level specification of critical system context can be done by an attacker based on the knowledge of the typical functionality of an ACC system and be applied to any ACC system with the same functional specification.
TABLE III: Partial safety context table for an ACC system.
@@ 4 RS>0 No
Figure 3: Optimization-based adversarial patch generation.
III-C2 Adaptive Adversarial Patch Generation
@@
Runtime Optimization-based Adversarial Patch Generation. To address challenge C2, we formulate the attack as the following runtime optimization problem:
min $\sum_{d\in{RD}_{t}}{-\triangledown g(d,\theta)}+\lambda||\Delta_{t}||_{p}$ (1)
s.t. $Patch_{t}=\Delta_{t}*M_{t}$ (2)
$Patch_{t}\in[\mu-\sigma,\mu+\sigma]$ (3)
$Area(Patch_{t})\subset BBox(LV)_{t}$ (4)
$X^{adv}_{t}=X_{t}+Patch_{t}$ (5)
$[{RD},RS]_{t}=LVD_{\theta}(X^{adv}_{t-1})$ (6)
$u_{t}=ACC(s_{t},[{RD},{RS}]_{t})$ (7)
@@
where Eq. 1 is an objective function that aims at accelerating the Ego vehicle as soon as possible to cause a forward collision.Directly decreasing the probability of a lead vehicle or its bounding box (BBox) cannot change the ACC system behavior. We instead design an objective function that increases $RD_{t}$ as much as possible while keeping the perturbation value of the adversarial patch imperceptible to human eyes.
In Eq. 1, $g(d)$ is an approximate polynomial function of $d$ that fits the trend of the trajectory of the relative distance ${RD}_{t}$, predicted by the DNN model with weight parameters $\theta$. "-" is a negative sign that converts our goal of maximizing the relative distance to minimizing the proposed objective function. For example, when the gradient of the relative distance trajectory, $g(d)$, is negative, minimizing "$-\triangledown g(d)$" will slow down the decrease of $g(d)$ and is equivalent to maximizing the relative distance. We adopt the gradient of $g(d)$ in the objective function instead of using ${RD}_{t}$ itself in order to avoid sharp changes in the predicted relative distance value, which might be easily detected by some anomaly detection mechanisms.
In Eq. 5, the perturbation is added to the original image input $X_{t}\in\mathbb{R}^{H\times W\times C}$ in the form of an adversarial patch $Patch_{t}\in\mathbb{R}^{H\times W\times C}$, represented as a matrix of pixels with height H, width W, and C color channels. $\lambda$ is the weight parameter of the p-norm regularization term, designed to minimize the perturbation value of the patch for stealthiness. We limit the perturbation value within the Kalman filter noise parameters ($\mu$,$\sigma$) (Eq. 3), which ensures the perturbation is not corrected by the sensor fusion. We also constrain the adversarial patch inside the BBox of the lead vehicle (Eq. 4) to enhance attack effectiveness, minimize the perturbation area for stealthiness, and reduce computational cost.
@@
Primary Attribution Detection and Patch Update. As mentioned in Section III-B, a major challenge (C3) in the design of runtime attacks is the changes in the size and location of the lead vehicle in the perceived image frames. To address this challenge, the attacker needs to update the generated patch dynamically according to the approximate DNN outputs. In this paper, we adopt an object detection method [71] to detect and track the real-time location and size of the lead vehicle and concentrate the attack perturbation within the detected BBox of the lead vehicle. In production ACC with object detection features [21], given proper access, the attack can skip this step and use the stock prediction results.
@@
After getting the BBox, we utilize a primary attribution algorithm [83] to quantify the relationship between input features and output predictions. Through this exploration, we try to identify the important pixels within the BBox of the lead vehicle that contribute the most to the predictions of $RD_{t}$. The input pixels with high weights identified by the attribution algorithm are marked by unit value in the mask matrix $M_{t}$, and the remaining pixels are assigned zero values. This mask matrix is then multiplied by the perturbation $\Delta_{t}$ to generate the adversarial $Patch_{t}$ (in Eq. 2). This step is useful as it can filter out non-important pixels in the inputs to reduce the number of perturbed pixels to improve the effectiveness of optimization-based attacks and reduce computation costs of runtime attacks.
@@
Finally, we develop a new initialization algorithm to shift the patch position and adjust its size when the detected BBox changes (Eq. 9-11). We shift the attack vector toward the new position of the detected BBox of the lead vehicle with a magnitude of ($x_{t}-x_{t-1}$, $y_{t}-y_{t-1}$), where ($x_{t-1}$, $y_{t-1}$) and ($x_{t}$, $y_{t}$) are the centers of BBox at previous and current control cycles, respectively (Eq. 9). We then expand the adversarial patch attack vector ($Patch$) to the dimension that matches the size of the newly detected BBox of the lead vehicle. Instead of reinitializing the whole attack vector matrix with random or zero values, which will reset the whole optimization process, we keep the previous patch values and intermediate variables and only initialize the newly expended units (Eq. 10-11). An example is shown in Fig. 5.
This algorithm maintains a continuous optimization process across two consecutive perception cycles, which is critical in satisfying the real-time constraints. Fig. 5 shows a visualization of how the adversarial patch affects the DNN predictions.
IV Safety Intervention Simulation
@@
To evaluate the safety of DNN-based ACC systems under attacks, we enhance the default OpenPilot and CARLA simulation platform (see Section II-C) to be more representative of real-world ADAS, by developing a safety intervention simulator and mechanisms for priority-based dispatching of control commands to CARLA and fusion of camera and radar data (see Appendix A). An overview of the simulation platform is shown in Fig. 6 (with the orange parts representing our new implementations) and presented next.
@@
To fill the gap in considering safety interventions and address the challenge of ensuring the combination of human driver and vehicle safe (see Section II-B), we implement and integrate three levels of safety interventions in the OpenPilot software (see Fig. 6), including ADAS safety features (e.g., AEB and FCW), basic car safety constraint checking on control commands, and driver interventions.
@@
The AEBS processes inputs from LVD outputs (after sensor fusion), including relative distance ($RD$) and relative speed ($RS$), and current speed $V_{Ego}$ (see Fig. 1). The average driver reaction time ($T_{react}$) is set to the commonly used constant value of 2.5s [97, 76]. Time thresholds, namely $ttc$ (time to collision), $t_{fcw}$ (FCW time), $t_{pb1}$ (1st phase partial brake time), $t_{pb2}$ (second phase partial brake time), and $t_{fb}$ (full brake time), are then calculated as follows:
$ttc$ $=RD/RS$ (12)
@@
As shown int Fig. 8, when $ttc$ falls below $t_{fcw}$, $t_{pb1}$, $t_{pb1}$, and $t_{pb1}$, a corresponding action (warning or brake with 90%, 95%, 100% force) is executed. Applying the brake value blocks other controls from the ADAS. More details of AEBS design and testing are provided in Appendix IX.
In reality, when OpenPilot is installed on a car model, some models lose the AEBS functionality [9], while others retain it. Also, AEBS might rely on a separate ADAS camera [82], distinct from the OpenPilot camera, and a potential compromise of the AEBS camera data is possible. Thus, we consider three scenarios for AEBS interventions: (1) AEBS enabled, and AEBS camera data is uncompromised; (2) AEBS enabled, but AEBS camera data is compromised; and (3) AEBS is disabled (see Section V-C2 and Table VII).
Safety Constraint Checker. The OpenPilot safety mechanisms are implemented in its control software and the Panda CAN interface. Panda is a universal OBD adapter developed by Comma.ai [7] that provides access to almost all car sensors through the CAN bus and also enforces some safety constraints over output commands. However, when integrated with the CARLA driving simulator, OpenPilot does not utilize Panda software or hardware; thus Panda safety checks are inactive.
[…]
Driver Reaction Simulator. To assess driver interventions, we develop a driver reaction simulator. The simulated driver is notified when any safety alerts are raised by the ADAS (e.g., FCW) or when the driver observes any abnormalities in the vehicle’s status or camera user interface (UI). These ACC abnormalities include hard braking, unexpected acceleration, the vehicle’s speed exceeding the cruising speed by more than 10%, or the mean perturbation value in the UI surpassing a noticeable threshold, with a default value of 15% representing an alert driver ($Patch.mean()>0.15$). We assume a very alert driver who can notice any anomalies that occur within a single control cycle (10ms). A predefined emergency response will be issued by the driver accordingly (see Table IV), taking effect 2.5 seconds later (average driver reaction time). To mimic the human driver’s braking behavior, we adopt a braking curve function from previous research [97].
TABLE IV: Driver simulator: activation conditions and reactions.
@@
Priority-based Control Command Dispatcher. With multiple safety mechanisms in place, there might be conflicts among the control commands issued by the OpenPilot ACC controller and those generated by the safety interventions. To resolve such conflicts, we design a command dispatcher to transmit output control commands to the CARLA actuators from various sources (e.g., ACC, AEB, simulated driver) based on their priorities, with high-priority commands overwriting low-priority ones (see Fig. 6 and Fig. 8). The simulated driver’s actions have a higher priority than regular ACC outputs, and control actions from the AEB have the highest priority. The driver’s actions will be executed 2.5s (average driver reaction time) after safety alerts (e.g., FCW) or noticing other ACC malfunctions (see Table IV). ACC commands will be blocked or disengaged when AEB or driver interventions are triggered.
@@
V Evaluation in Simulated Environment
This section presents the evaluation of our proposed context-aware attack strategy (referred to as CA-Opt) using the enhanced simulation platform presented in Section IV.
V-A Methodology
We study the following research questions by comparing the effectiveness of CA-Opt attack to several baseline attack methods in causing safety hazards (Section V-B) and evading different safety interventions (Section V-C):
@@ V-A Methodology
RQ3: Does a perception input attack achieve better performance than direct perception and control output attacks?
Baselines. We design three baseline attack strategies to answer these questions (see Table V).
TABLE V: Overview of proposed and baseline attack strategies.
@@
CA-APGD Context-Aware Context-Aware AutoPGD 1000
@@
To assess the effectiveness of our optimization-based adversarial patch method in strategically selecting the attack values, we compare it to two baselines: CA-Random, which introduces random perturbations to perception inputs, and CA-APGD, which uses a state-of-the-art gradient-based method, Auto-PGD [31], to determine the perturbation values. Since the original Auto-PGD is designed for misclassification, which does not work for attacking ACC, we change the goal function (to $\triangledown g(d)$, see Eq. 1) to maximize relative distance prediction. We also limit the iteration number to 5, the maximum number of control cycles (100Hz) within a perception cycle (20Hz)). Both baselines use the same context-aware method as the proposed CA-Opt attack strategy to choose the attack start time and duration. To evaluate our method’s efficiency in selecting the timing and duration of attacks, we design another baseline (Random-Opt) that selects a random start time uniformly distributed within [5, 40] seconds and a random attack duration uniformly distributed within [0.5, 2.5] seconds. To concentrate only on the effect of different start times and durations, Random-Opt shares the same adversarial patch generation method as CA-Opt. For a fair comparison, we confine perturbations to the detected bounding box (BBox) of the lead vehicle and update the BBox size and position with our proposed patch updating algorithm (Section 5).
@@
Our experiments are done on Ubuntu 20.04 LTS, with OpenPilot v0.8.9 and CARLA v9.11. A single simulation of OpenPilot contains 5,000 time steps, and each step lasts about 10 ms, which equals 50 seconds in total. However, if an attack leads to a collision, the simulation ends earlier. For Random-Opt, we randomly select ten start times and five durations for each test scenario and repeat them five times, which results in 1,000 simulations. The same total number of simulations is done for CA-Opt and other baselines.
V-B Attack Success Rate in Causing Hazards